OpenStudy (anonymous):

how do you solve: x^2 + y^2+8x-16y+44?

OpenStudy (amistre64):

by making a circle of an elipse out of it

OpenStudy (amistre64):

looks cicley to me

OpenStudy (anonymous):

your no help

OpenStudy (amistre64):

(x^2 +8x +16) + (y^2 -16y +64) = -44+16 + 64 (x+4)^2 + (y-8)^2 = 36

OpenStudy (amistre64):

im not?

OpenStudy (amistre64):

the "solution" is a circle of radius 6 centered at (-4,8), but what do I know lol